One of the angles of a triangle is equal to the sum of the other two angles. If the ratio of the other two angles is 4:5, find the angles of the triangle.
Answer by RAY100(1637)

Let 1st angle = 4x,,,,,,and 2nd angle = 5x
.
Then 3rd angle is the sum of the first 2 or ,,,,4x+5x = 9x
.
remember that the sum of the angles of a triangle is 180
.
4x + 5x + 9x + 180
.
18 x = 180
.
x=10
.
The 1st angle = 4x = 40,,,2nd angle = 5x =50, and 3rd angle =9x = 90
.
check,,,,40 +50 +90 = 180 ok, ratio of ist 2 angles is 40/50 = 4/5,,ok